A rigid flange coupling is used to connect a 45KW 1440rpm electric motor to a centrifugal pump. The starting torque of the motor is 225% of the rated torque. There are 8 bolts and their pitch circle diameter is 150mm . The bolts are made of steel 45C8 (sigma yt 380N/mm2 ) the factor of safety is 2.5 . Determine the diameter of the bolts. Assume (sigma sy 0.577 sigma yt) . Assume that the bolts are finger tight is reamed and ground holes.

1
Expert's answer
2021-11-19T00:07:30-0500

Permissible shear stress,"\tau=\frac{s_{Sy}}{FS}"

"=\frac{0.577s_{st}}{2.5}=87.7N/mm^2"

Direct shear stress in bolt

"P_1'=P_2'=\frac{P}{Number\ of\ bolts}"

"=\frac{45*10^3}{8}=5625N"

Shear area of 8 bolts="8(\frac{πd^2}{4})mm^2"

P="8(\frac{πd^2}{4})*\tau"

"d^2=\frac{45*10^3*4}{8*π*87.7}"

"d^2=81.664"

"d=9.03mm"

Hence standard size of bolt is M10